Outline

• ATM Adaptation Layer • ATM Signaling

• An ATM Signaling Protocol

ATM Cell

GFC (4b), VPI (8b), VCI (16b), Type (3b), CLP (1b), Header CRC (8b), Data (384b/48B).

Correction from last lecture, where VCI was inadvertently given as 12 bits long.

ATM AAL

An ATM payload is too small to carry much useful data. Therefore, part of the ATM definition is a layer to do segmentation and reassembly. This layer is called ATM Adaptation Layer or AAL.

Different AALs serve different functions:

• AAL 1 and AAL 2 support guaranteed bit rates (voice).

• AAL 3/4 and AAL 5 support packet data.

The packets given to and returned by the AAL are called Protocol Data Units (PDUs).

AAL 3/4

Per-PDU header (CPI/1, Btag/1, BASize/2), 1 - (2^{16} − 1) bytes of data, trailer (Pad/1-4, Etag/1, Length/2). Overhead is 8 bytes per PDU.

Per-cell header (Type/2b, Seq/4b, MuxID/10b) and trailer (Length/6b, CRC-10/10b) leave up to 44 bytes for data. Overhead is 4 bytes per cell.

Total PDU overhead is 8 + 4 × \frac{n}{44}. This leaves maximum efficiency below 83%, which is low. Also, computing one CRC every 44 bytes of data can be slow if done in software (the header CRC is likely computed in hardware, since it’s AAL-independent).

AAL 5

AAL 5 was introduced with the purpose of lowering the overhead of AAL 3/4. AAL 5 uses the “user signaling bit” in the ATM header to mark the end of a frame.

Format is: Data/1-(2^{16} − 1), Pad/0-47, Reserved/2, Length/2, CRC-32/4.

Total PDU overhead is 8 bytes, and there is only one data CRC computation per per PDU.

ATM Signaling

The purpose of signaling is the establishment, maintenance, and teardown of connections.

In ATM, a host or switch sends a request towards a destination. The request is routed to the destination in the same way that a connectionless datagram would be routed. Each switch along the route inspects the request, and may modify it or allow it to have a side effect. At the end, the destination returns a response, which is likewise inspected by every switch along the route.

One of the side effects of a request could be setting up a virtual channel, for example.

The above and what follows are not going to be tested on... they are not in the book or anything.

6-1

SPANS

SPANS (Simple Protocol for ATM Network Signaling) is a non-standard protocol developed at Fore Systems in 1992.

Operations supported include setting up and tearing down a unidirectional (unicast or multicast) virtual channel. There are no maintenance operations.

A multicast virtual channel is added one end-point at a time, and the connection request has to come from the sender (no receiver-initiated join). The switches support multicast in hardware.

Connection Setup

Addresses are of the form (switch ID, port number).

A host sends a connection request to its attached switch. The request contains a destination address. The switch knows how to route to all other switches in the network, so forwards the request towards the destination. The response must follow the inverse path.

QOS Negotiation

A connection request contains a QOS (quality of service) which is defined as average bandwidth, peak bandwidth, and burst size. If a switch is unable to allocate the QOS, it returns failure (without going all the way to the destination).

A request actually carries both a minimum and a desired QOS. Switches and the destination are allowed to reduce the desired QOS, but if they would have to reduce it below the minimum QOS, they must fail the request instead.

Link state, configuration

SPANS also defines “hello” packets to be exchanged on a link for the purpose of verifying that the link is active. The packets carry the address of the sender and the port number (if any) sent on. This information allows autoconfiguration of hosts, since the pair (switch ID, port ID) is the host address.

These same packets also provide enough information to gather the information needed for a link-state routing protocol.

Finally, each link has a predefined VCI for signalling, and a predefined VCI for broadcasting.

SPANS NNI

The “Network-to-Network Interface” signaling protocol allows for:

• setting up virtual paths as well as virtual channels.

• distributing link-state information.

• broadcasting (used for ARP).

Summary

ATM is designed for high speed: fixed cell size and simple cell processing mean that cells can be processed in hardware.

ATM is designed for both local networks (high speed) and wide-area networks (point-to-point links).

ATM is designed to take over the world (telephone companies and computer networks). Will it?